OpenStudy (mony01):

Find solutions of the given IVP. (1+e^t)dy/dt+e^(t)y=0 y(0)=1

ganeshie8 (ganeshie8):

hint : \(\large \left[(1+e^t)y\right]'\)

OpenStudy (mony01):

what

ganeshie8 (ganeshie8):

remembe the product rule ?

OpenStudy (mony01):

yea

ganeshie8 (ganeshie8):

\(\large (f*g)' = ?\)

OpenStudy (mony01):

f'g+fg'

ganeshie8 (ganeshie8):

expand the derivative of product in hint above, what do you get ?

OpenStudy (mony01):

e^t(y')+(1+e^t)y

ganeshie8 (ganeshie8):

try again

OpenStudy (mony01):

i dont see a mistake

ganeshie8 (ganeshie8):

sure ?

ganeshie8 (ganeshie8):

it seems you're doing f'g' + fg

ganeshie8 (ganeshie8):

instead of fg' + f'g

OpenStudy (mony01):

oh e^t(y)+(1+e^t)(y')

ganeshie8 (ganeshie8):

yes! isn't that same as the left hand side of your differential equation ?

OpenStudy (mony01):

yes

ganeshie8 (ganeshie8):

\[\large (1+e^t)dy/dt+e^ty=0 \] using product rule, becomes : \[\large \left[(1+e^t)y\right]' = 0\]

ganeshie8 (ganeshie8):

which function has "0" as derivative ?

OpenStudy (mony01):

y?

ganeshie8 (ganeshie8):

derivative of a constant function is 0

ganeshie8 (ganeshie8):

because a constant function never changes

ganeshie8 (ganeshie8):

\[\large \left[(1+e^t)y\right]' = 0\] or you could integrate both sides so that the derivative disappears on left hand side : \[\large \int \left[(1+e^t)y\right]' dt = \int 0 ~dt\] \[\large (1+e^t)y = C\]

ganeshie8 (ganeshie8):

plugin the initial condition and solve C

OpenStudy (mony01):

plug in 1 for y

ganeshie8 (ganeshie8):

yes 1 for y and 0 for t

OpenStudy (mony01):

c=2

ganeshie8 (ganeshie8):

yes plug that back into the final equation : \[\large (1+e^t)y = 2\]

ganeshie8 (ganeshie8):

or \(\large y = \dfrac{2}{1+e^t}\) is the solution function for the differential equation

OpenStudy (mony01):

thanks

OpenStudy (mony01):

how would i know which part of the equation to equal it to 0?

ganeshie8 (ganeshie8):

what variables do you see in the given equation ?

OpenStudy (mony01):

t,y', and y

ganeshie8 (ganeshie8):

thats true, lets keep y' aside. `t` and `y` are the variables and you're trying to solve for the function `y` right ?

ganeshie8 (ganeshie8):

and `y` is a function of `t`

ganeshie8 (ganeshie8):

y(0) = 1 means when t = 0, the value of y is 1

ganeshie8 (ganeshie8):

\(\large \large y = \dfrac{2}{1+e^t}\) assumption is that y is a function of t : \(\large y(t) = \dfrac{2}{1+e^t}\)

OpenStudy (anonymous):

I think of it as initial time idk if that'll help though.

OpenStudy (mony01):

so get t and y' together to equal 0

ganeshie8 (ganeshie8):

y(0)=1 you're asking how to interpret this, right ?

OpenStudy (mony01):

im asking why you chose (1+e^t)y'=0 and not e^t y

ganeshie8 (ganeshie8):

when did we choose that ?

ganeshie8 (ganeshie8):

we never set (1+e^t)y'=0

OpenStudy (mony01):

at the beginning using the product rule

ganeshie8 (ganeshie8):

Your starting equation : \(\large (1+e^t)dy/dt+e^{t}y=0 \) \(\large (1+e^t)y'+e^{t}y=0 \)

ganeshie8 (ganeshie8):

now wrap that using the product rule in reverse

ganeshie8 (ganeshie8):

entire left hand side can be compressed as : \(\large \left[(1+e^t)y\right]'\)

ganeshie8 (ganeshie8):

yes ?

OpenStudy (mony01):

why? where did e^t y go

ganeshie8 (ganeshie8):

the derivative is NOT just to the y, it was to the entire expression inside square parenthesis

ganeshie8 (ganeshie8):

it did not go anywhere, expand it using product rule

OpenStudy (mony01):

i thought product rule was just for (1+e^t) and y'

ganeshie8 (ganeshie8):

\(\large \large \left[(1+e^t)y\right]'\) is exactly same as : \(\large (1+e^t)y'+e^{t}y\)

OpenStudy (mony01):

oh i didn't know that

ganeshie8 (ganeshie8):

work it using product rule and convinces yourself they are same

ganeshie8 (ganeshie8):

\(\large \large \left[(1+e^t)y\right]' = ?\)

OpenStudy (mony01):

it equals the same thing, so how can I work backwards to obtain that expression?

ganeshie8 (ganeshie8):

you need to *see* it and think in reverse

OpenStudy (mony01):

oh thats kinda hard

ganeshie8 (ganeshie8):

\(\large \large (1+e^t)y'+e^{t}y\) there is no rule for writing it as derivative of product without thinking in reverse

ganeshie8 (ganeshie8):

however you can convince urself more that it can be written asderivative of product like this : \(\large (1+e^t)y'+e^{t}y\) derivative of the first factor in first term \((1+e^t) \) comes as first factor in second term. So thats a huge hint that you can wrap it as : \(\large (fg)'\)

OpenStudy (mony01):

so in all of these kinds of problems I need to do the same thing, like think about what expression would equal that when using the product rule

ganeshie8 (ganeshie8):

also notice that one of the functions is always \(\large y\). its not hard at all after you practice 2-3 problems

ganeshie8 (ganeshie8):

if you want a practice example, try to write below as derivative of product : \(\large x^2y' + 2xy\)

OpenStudy (mony01):

[(x^2)y]'

ganeshie8 (ganeshie8):

perfect ! want another example ?

OpenStudy (mony01):

so for dy/dt+2ty=t^3

ganeshie8 (ganeshie8):

that won't work this simple

ganeshie8 (ganeshie8):

as left side cannot be written as derivative of product

ganeshie8 (ganeshie8):

the good news is we can force it to be written as derivative of product

ganeshie8 (ganeshie8):

multiply something bothsides such that the lefthand side becomes derivative of product

ganeshie8 (ganeshie8):

there is a procedure for finidng that "something"

ganeshie8 (ganeshie8):

heard of "integrating factor" before ?

OpenStudy (mony01):

yea i think

ganeshie8 (ganeshie8):

what do you know about it ?

OpenStudy (mony01):

is it using this \[\mu=e ^{\int\limits_{}^{}p(t)dt}\]

ganeshie8 (ganeshie8):

exactly! find it

ganeshie8 (ganeshie8):

and multiply that both sides of the equation

OpenStudy (mony01):

so do I solve for y or c

OpenStudy (mony01):

okay i got y=1/2t^2 - 1/2 +c

ganeshie8 (ganeshie8):

whats your integrating factor ?

ganeshie8 (ganeshie8):

dy/dt+2ty=t^3

ganeshie8 (ganeshie8):

integrating factor = \(\large e^{\int 2t~dt} = e^{t^2}\) right ?

OpenStudy (mony01):

yea

ganeshie8 (ganeshie8):

multiply that through out the equation

ganeshie8 (ganeshie8):

\(\large dy/dt+2ty=t^3\) multiiply the integrating factor \(\large \color{gray}{e^{t^2}}\) through out, you get : \(\large \color{gray}{e^{t^2}}dy/dt+\color{gray}{e^{t^2}}2ty=\color{gray}{e^{t^2}}t^3\)

ganeshie8 (ganeshie8):

now, you can blindly write left hand side as \(\large \left[y \color{gray}{e^{t^2} }\right]' \)'

OpenStudy (mony01):

yes

ganeshie8 (ganeshie8):

\(\large \left[y \color{gray}{e^{t^2} }\right]' =\color{gray}{e^{t^2}}t^3 \)

ganeshie8 (ganeshie8):

integrate both sides

ganeshie8 (ganeshie8):

\[\large \int \left[y \color{gray}{e^{t^2} }\right]' ~dt = \int ~\color{gray}{e^{t^2}}t^3~dt \]

ganeshie8 (ganeshie8):

left hand side, integral and derivative eat eachother out giving you : \[\large y \color{gray}{e^{t^2} } = \int ~\color{gray}{e^{t^2}}t^3~dt \]

ganeshie8 (ganeshie8):

you will have to evaluate the right hand side integral

OpenStudy (mony01):

i know the integration but how can i find it by hand, product rule?